A 2.00-kg toy accelerates from rest to 3.00 m/s in 8.00 s on rough surface of uk=0.300. Find the applied force f



1
Expert's answer
2021-03-03T11:24:34-0500

Let's first find the acceleration of the toy:


"a=\dfrac{v-v_0}{t}=\dfrac{3.0\ \dfrac{m}{s}-0}{8.0\ s}=0.375\ \dfrac{m}{s^2}."

Let's apply the Newton's Second Law of Motion:


"F_{appl}-\mu_kmg=ma,""F_{appl}=m(a+\mu_kg),""F_{appl}=2.0\ kg\cdot(0.375\ \dfrac{m}{s^2}+0.3\cdot9.8\ \dfrac{m}{s^2})=6.63\ N."